Suppose that Jamess income in Year 1 was $43,000. Over the next four years his income increases by 5% per year. At the same time. the economy experiences an inflation rate of 4% per year. a. Complete the table below. Do not round your intermedlate calculations. Enter your final responses as a whole number for Nominal Income and Real Income Enter your final responses for the price Index rounded to one declmal place. Year Nominal incomePrice index Real income $4300 $43,000 100 2 3 4 5 b. At the end of Year 5, Jamess real income will be$
Answer #1

(a)

Year 2

Nominal income increases by 5% per year.

Nominal income in Year 1 = $43,000

Nominal income in Year 2 = $43,000 + ($43,000 * 0.05) = $43,000 + $2,150 = $45,150

Inflation rate is 4% per year.

Price index in Year 1 = 100

Price index in Year 2 = 100 + (100 * 0.04) = 100 + 4 = 104

Real income in Year 2 = Nominal income in year 2 * (Price index in Year 1/Price index in Year 2)

Real income in Year 2 = $45,150 * (100/104) = $43,413

Year 3

Nominal income in Year 2 = $45,150

Nominal income in Year 3 = $45,150 + ($45,150 * 0.05) = $45,150 + $2,257 = $47,407

Price index in Year 2 = 104

Price index in Year 3 = 104 + (104 * 0.04) = 104 + 4.16 = 108.16 or 108.2

Real income in Year 3 = Nominal income in year 3 * (Price index in Year 1/Price index in Year 3)

Real income in Year 3 = $47,407 * (100/108.2) = $43,814

Year 4

Nominal income in Year 3 = $47,407

Nominal income in Year 4 = $47,407 + ($47,407 * 0.05) = $47,407 + $2,370 = $49,777

Price index in Year 3 = 108.2

Price index in Year 4 = 108.2 + (108.2 * 0.04) = 108.2 + 4.328 = 112.528 or 112.5

Real income in Year 4 = Nominal income in year 4 * (Price index in Year 1/Price index in Year 4)

Real income in Year 4 = $49,777 * (100/112.5) = $44,246

Year 5

Nominal income in Year 4 = $49,777

Nominal income in Year 5 = $49,777 + ($49,777 * 0.05) = $49,777 + $2,489 = $52,266

Price index in Year 4 = 112.5

Price index in Year 5 = 112.5 + (112.5 * 0.04) = 112.5 + 4.5 = 117

Real income in Year 5 = Nominal income in year 5 * (Price index in Year 1/Price index in Year 5)

Real income in Year 5 = $52,266 * (100/117) = $44,672

Following is the complete table -

Year Nominal Income Price Index Real Income
1 $43,000 100 $43,000
2 $45,150 104 $43,413
3 $47,407 108.2 $43,814
4 $49,777 112.5 $44,246
5 $52,266 117 $44,672

(b)

At the end of Year 5, James's real income will be $44,672.